64. A common mistake when solving equations is the following:
The equation: 2(x- 2) = x+ 3
First step in solving: 2x- 2= x + 3
Write a clear explanation of what error has been made. What could be done to
avoid this error?

Answer by stanbon(48558)

A common mistake when solving equations is the following:
The equation: 2(x- 2) = x+ 3
First step in solving: 2x- 2= x + 3
MISTAKE: 2(x-2) does not equal 2x-2; it equals 2x-4
Write a clear explanation of what error has been made. What could be done to
avoid this error?
Use an explicit application of the distributive law:
2(x-2)=2*x-2*2=2x-4
=============
Cheers,
Stan H.